OpenStudy (anonymous):

O(n) is an odd-number function. Find O(13), where O(n) = 2n - 1.

hartnn (hartnn):

n=13 2*13-1 =.. ?

OpenStudy (anonymous):

what??????

OpenStudy (anonymous):

O(13)=2(13)-1=25 that's it!!!